# FLUXSynID: A Framework for Identity-Controlled Synthetic Face Generation with Document and Live Images

### v2

@@ 4.3 PuLID-based Image Generation
PuLID (Pure and Lightning ID Customization) [22] personalizes text-to-image diffusion models using identity-specific embeddings derived from reference images. Unlike traditional fine-tuning methods, PuLID allows for rapid identity injection without the need for retraining each subject. It uses a dual-path training strategy: one path conditioned on text only, and the other on both text and identity embeddings, with contrastive and identity loss to align outputs.
Within FLUXSynID, we employ PuLID-FLUX-v0.9.1 [54], an adaptation of PuLID integrated with the pre-trained and frozen FLUX.1 [dev] diffusion model [3]. As shown in Figure 3, the synthetic document image $I_{D}$ and the LivePortrait-modified image $I_{LL}$ are passed through PuLID’s identity encoder. The resulting embeddings are averaged to form the identity condition $c_{ID}$, capturing both static identity features and dynamic pose or expression cues.
PuLID introduces dedicated cross-attention layers into FLUX.1 [dev]’s transformer blocks to integrate $c_{ID}$. Unlike text embeddings, which are fused via global modulation and token concatenation, identity embeddings interact directly with the FLUX.1 [dev] transformer layers via these specialized attention modules.
Simultaneously, the original identity prompt $P$ is transformed into a contextual prompt $P^{\prime}$, describing the same individual in an airport setting, and encoded via CLIP-L [46] and T5-XXL [47] to produce the textual condition $c_{P^{\prime}}$.
In parallel, edges from $I_{LL}$ are extracted and encoded into a latent representation $z_{t,\text{edges}}$ using the pre-trained FLUX.1 [dev] VAE encoder ($E_{\text{VAE}}$[3]. Simultaneously, a Gaussian noise latent image $z_{t,\text{noise}}$ is generated with the same seed as the document image. Both latent images are concatenated across the channel dimension to form $z_{t}^{*}$:
$z_{t}^{*}=z_{t,\text{edges}}\mathbin{\|}z_{t,\text{noise}}.$ (2)
The concatenated latent image $z_{t}^{*}$ serves as the input for the FLUX.1 [dev] diffusion model equipped with a specialized FLUX.1 [dev] Canny LoRA adapter [3]. This LoRA adapter modifies the input layer of FLUX.1 [dev] and is explicitly trained for edge-based image-to-image conditioning.
Therefore, the final generation is conditioned on three signals: identity embedding ($c_{ID}$), modified text embedding ($c_{P^{\prime}}$), and latent edge representation ($z_{t,\text{edges}}$), enabling realistic, identity-consistent live capture synthesis. The conditional denoising process in PuLID-based live image generation is formulated as:
@@
where $\epsilon_{\theta^{Canny}}$ represents FLUX.1 [dev] model combined with the FLUX.1 [dev] Canny LoRA adapter [3].
@@ 5.1 Implementation Details
Prompt Generation.
Identity prompts were generated using the Qwen2.5 [57] model, based on sampled identity attributes. Attribute inclusion probabilities and conflict rules were applied to generate semantically rich prompts optimized for conditioning FLUX.1 [dev] [3].
LoRA Training for Document Image Generation.
A LoRA adapter was trained using FluxGym [9] on 830 images from the Chicago Face Database [37, 38, 32], with rank $r=16$, learning rate 0.0002, and 18 training epochs. Image captions were generated using the InternVL2.5 [8] vision-language model. The LoRA module was activated during inference via a unique token (“s7ll2f4h”), and additional parameters in the CLIP-L [46] encoder were also fine-tuned.
Document Image Generation.
Document images were synthesized through the ComfyUI-based [10] implementation of FLUX.1 [dev] [3], conditioned on Qwen2.5 prompts. Guidance scales were sampled randomly between 1.7 and 2.5 to balance fidelity with identity diversity. Each image was generated using $t=20$ diffusion steps.
Live Image Generation.
LivePortrait [21] and PuLID-FLUX-v0.9.1 [54] were used via modified ComfyUI-based [10] implementations. A fixed guidance scale of 4 and $t=20$ steps were used with PuLID, as recommended by the authors. Arc2Face [42] was run via the official implementation with default hyperparameters.
Dataset Composition.
A total of 15,000 synthetic identities were generated across 14 attribute classes: gender, age, region of origin, body type, eye shape, lip shape, nose shape, face shape, hairstyle, hair color, eyewear, facial hair, skin type, and ICAO-compliant headwear [56]. The gender, age, and region of origin attributes were included in 100% of prompts to guarantee consistent representation. The probabilities for age, region of origin, and body type attributes were uniformly distributed, whereas other attributes were subjectively assigned, reflecting author preference instead of real-world distributions. Researchers are encouraged to define custom distributions for task-specific dataset generation.
Of the 15,000 identities, 111 were excluded due to Arc2Face [42] failing to extract identity embeddings. These identities were removed to ensure each identity includes outputs from all three live image generation methods.
All images were generated at a resolution of $1024\times 1024$, except for Arc2Face-based live images, which were constrained to $512\times 512$ due to implementation limitations.
@@ 5.2 Similarity-Based Identity Filtering
Due to prompt similarity and inherent biases in the FLUX.1 [dev] [3] model, some synthetic identities may appear overly similar. Unlike ONOT [14], which enforces strict identity separation by removing all similar identities detected by a face recognition system (FRS), FLUXSynID allows a small degree of similarity, reflecting real-world datasets where a limited False Match Rate (FMR) is acceptable. For instance, FRONTEX guidelines [18] permit an FMR of 0.1%. As dataset size grows, the probability that a randomly selected identity $i$ falsely matches at least one other identity in the dataset is:
@@
To manage this behavior, FLUXSynID employs iterative similarity-based filtering. Starting from an initial set of document images $D_{o}$, pairs of identities with cosine similarity exceeding a threshold are identified. The filtering process iteratively removes identities involved in the highest number of such matches. This continues until the dataset-wide FMR reaches or falls below a predefined static FMR target:
@@
Experiments in this study leverage two open-source face recognition models: ArcFace [12] and AdaFace [28]. For each model, two decision thresholds are selected, corresponding to FMR targets of 0.1% and 0.01%. These thresholds are derived from 340k impostor verification attempts on identities from the CFD [37, 38, 32] dataset, yielding values of 0.423 and 0.497 for ArcFace, and 0.253 and 0.334 for AdaFace, respectively. Accordingly, the $\text{FMR}_{\text{target}}$ in Equation (5) is set to 0.1% or 0.01%, depending on the chosen threshold.
@@
Table 1: Remaining identities after similarity-based filtering at different FMR targets using ArcFace [12] and AdaFace [28]. Percentages are relative to the initial 14,889 identities.
@@ 5.3 Evaluation
Visual Results.
Figure 4 shows sample identities from FLUXSynID, each consisting of a document image ($I_{D}$) and live capture images generated by LivePortrait ($I_{LL}$[21], PuLID ($I_{LP}$[22], and Arc2Face ($I_{LA}$[42]. Document images are frontal, feature neutral expressions, and have uniform white backgrounds, reflecting stylistic control via CFD-based LoRA fine-tuning.
Each live image generation method introduces different levels of variation. LivePortrait introduces subtle, controlled changes in pose and expression while maintaining consistent lighting and background, although minor blurring is sometimes evident. PuLID leverages text, edge maps, and identity embeddings to alter both background and pose, although it may occasionally cause identity drift. Arc2Face generates the most diverse outputs, introducing substantial changes in setting, lighting, and expression that closely mimic in-the-wild photographs.
@@
Prompt-Image Consistency.
Figure 5 shows synthetic document images generated from randomly sampled identity attributes. Generally, the images align well with the specified attributes, although some limitations are evident. Age estimation occasionally deviates, producing individuals who appear significantly younger or older than intended, consistent with findings in [41], which highlight FLUX.1 [dev] [3]’s difficulty in accurately representing age.
Figure 5: Examples of randomly sampled identity attributes and their corresponding synthetic document images ($I_{D}$) generated using FLUXSynID framework.
@@
While increasing the guidance scale improves prompt adherence, excessively high values can introduce artifacts and reduce realism. Empirically, guidance scales between 1.7 and 2.5 yielded the best trade-off between fidelity and prompt consistency.
Impact of Similarity-Based Filtering on Attribute Distribution.
To evaluate the potential bias introduced by similarity-based filtering (Section 5.2), Figure 6 compares the distributions of gender, age, and region of origin attributes before and after filtering.
Female identities were observed to be removed more frequently, reducing their overall representation. This indicates that FLUX.1 [dev] [3] generates less diverse female faces, leading to higher similarity scores and more frequent exclusion. Conversely, male and non-binary identities increased proportionally, indicating greater internal variability.
@@
Distribution of Identity Embeddings.
Figure 7 visualizes identity embeddings extracted by ArcFace [12] and AdaFace [28] across real and synthetic datasets. Synthetic StyleGAN2 [27] images were generated using the pSp [48] encoder, producing random identities in frontal poses. For ONOT [14], ICAO-compliant images from Subset 1 (the largest identity set) were used. FLUXSynID identities were filtered using FMR@0.01% thresholds (Section 5.2) to remove highly similar identities. Only document images were considered to maintain consistency across datasets, all of which contain document-like image types.
@@
Figure 8: Distribution of cosine similarity scores for mated (blue) and non-mated (orange) identity pairs across real and synthetic datasets, using the AdaFace FRS with an FMR@0.01% threshold.
@@
Table 2: KL divergence between similarity score distributions of synthetic and real datasets for mated and non-mated identity pairs. Lower values indicate greater distributional similarity.
To further evaluate FLUXSynID, cosine similarity scores for mated pairs (document image and its corresponding live image) and non-mated pairs (document image compared to live images from 100 different identities) are visualized as histograms in Figure 8. Comparisons are made to real-world FRLL [11] and FRGC [43] datasets.
@@
These observations are quantified in Table 2, which reports KL divergence between synthetic and real datasets. FLUXSynID-LL yields the lowest mated-pair divergence from FRLL, supporting its close visual match and controlled variation. For non-mated pairs, FLUXSynID-LA achieves the lowest divergence, though its mated-pair scores are higher. Overall, FLUXSynID-LL offers the most balanced similarity to FRLL across both pair types.
For FRGC, FLUXSynID-LA records the lowest mated-pair KL divergence, although FLUXSynID-LP (PuLID) visually resembles FRGC images more closely. While its KL values are slightly higher than Arc2Face’s, they still indicate strong distributional similarity.
In contrast, ONOT exhibits high KL divergence across all comparisons, indicating that its similarity score distributions differ substantially from those of real datasets. This observation aligns with Figure 8, where ONOT’s mated and non-mated distributions are visibly shifted compared to real data.
@@ 6 Conclusion
We present FLUXSynID, a synthetic face dataset generation framework capable of producing high-resolution, identity-consistent document and live capture images with fine-grained control over identity attributes. Combining prompt-driven diffusion synthesis with LoRA fine-tuning [24] and three complementary live image generation methods, FLUXSynID facilitates both controlled and naturalistic variations across identity instances. Experimental results demonstrate that FLUXSynID aligns more closely with the distributional characteristics of real-world datasets than existing synthetic alternatives, while its embedding space diversity supports realistic biometric evaluation scenarios.
Despite its strengths, FLUXSynID does have some limitations. Attributes such as age and subtle facial features are not always represented reliably, and similarity-based filtering may unintentionally skew demographic distributions. In contrast to ONOT, which enforces strict ICAO compliance [56] at the cost of a 73% reduction in dataset size [14], FLUXSynID does not enforce ICAO constraints. This decision reflects the limitations of current compliance tools, which frequently involve opaque checks and can result in compounding errors. Instead, FLUXSynID prioritizes dataset size and embedding space diversity over strict ICAO adherence, enhancing its flexibility for a broader range of biometric research applications.
@@ 7 Acknowledgments
This research was funded by the European Union under the Horizon Europe programme, Grant Agreement No. 101121280 (EINSTEIN).

### v3

@@ 4.3 PuLID-based Image Generation
PuLID (Pure and Lightning ID Customization) [21] personalizes text-to-image diffusion models using identity-specific embeddings derived from reference images. Similar to Arc2Face, PuLID enables rapid identity injection without requiring network retraining for each subject.
In FLUXSynID, we employ PuLID-FLUX-v0.9.1 [52], an adaptation of PuLID integrated with the pre-trained and frozen FLUX.1 [dev] [3] diffusion model. As shown in Fig. 3, synthetic document images ($I_{D}$) and LivePortrait-modified images ($I_{LL}$) are passed through PuLID’s identity encoder to extract static and dynamic identity features, which are averaged into a conditioning signal $c_{ID}$.
PuLID introduces dedicated cross-attention layers into FLUX.1 [dev]’s transformer blocks to integrate $c_{ID}$. These cross-attention layers inject identity information directly into the model by enabling focused interactions between the identity representation and the image tokens. This contrasts with the integration of text embeddings, which are incorporated via global modulation and token concatenation.
Simultaneously, the original identity prompt $P$ is transformed into a contextual prompt $P^{\prime}$, describing the same individual in an airport setting, and encoded via CLIP-L [44] and T5-XXL [45] to produce the textual condition $c_{P^{\prime}}$. To enhance structural fidelity, edges are extracted from $I_{LL}$ to provide spatial guidance. These edges are encoded into a latent image $z_{t,\text{edges}}$ using FLUX.1 [dev] VAE encoder [3] and are then concatenated along the channel dimension with a Gaussian noise latent image $z_{t,\text{noise}}$ to form $z_{t}^{*}$.
The concatenated latent image $z_{t}^{*}$ serves as the input for the FLUX.1 [dev] diffusion model equipped with a specialized FLUX.1 [dev] Canny LoRA adapter [3]. This LoRA adapter modifies the input layer of FLUX.1 [dev] and is trained for edge-based image-to-image conditioning.
The final generation is conditioned on three signals: identity embedding ($c_{ID}$), modified text embedding ($c_{P^{\prime}}$), and latent edge representation ($z_{t,\text{edges}}$). This ensures identity consistency while introducing natural pose and expression variations. The denoising process is defined as:
@@
where $\epsilon_{\theta^{Canny}}$ represents FLUX.1 [dev] model combined with the FLUX.1 [dev] Canny LoRA adapter.
@@ 5.1 Implementation Details
Prompt Generation. Identity prompts were generated using the Qwen2.5 [55] LLM, based on sampled identity attributes. Attribute inclusion probabilities and conflict rules were applied to generate semantically rich prompts optimized for conditioning FLUX.1 [dev] [3].
LoRA Training for Document Image Generation. A LoRA adapter was trained using FluxGym [9] on 830 images from the Chicago Face Database (CFD) [35, 36, 31], which provides high-resolution, front-facing portraits with neutral expressions and consistent lighting. The adapter was trained with rank $r=16$, learning rate 0.0002, and 18 training epochs. Image captions were generated using the InternVL2.5 [8] vision-language model. The LoRA module was activated during inference via a unique token (“s7ll2f4h”), and additional parameters in the CLIP-L [44] encoder were also fine-tuned.
Document Image Generation. Document images were synthesized through the ComfyUI-based [10] implementation of FLUX.1 [dev] [3], conditioned on Qwen2.5 prompts. Guidance scales were sampled randomly between 1.7 and 2.5 to balance fidelity with identity diversity. Each image was generated using $t=20$ diffusion steps.
Live Image Generation. LivePortrait [20] and PuLID-FLUX-v0.9.1 [52] were used via modified ComfyUI-based [10] implementations. A fixed guidance scale of 4 and $t=20$ steps were used with PuLID, as recommended by the authors. Arc2Face [40] was run via the official implementation with default hyperparameters.
Dataset Composition. A total of 15,000 synthetic identities were generated across 14 attribute classes: gender, age, region of origin, body type, eye shape, lip shape, nose shape, face shape, hairstyle, hair color, eyewear, facial hair, skin type, and ICAO-compliant headwear [25]. The gender, age, and region of origin attributes were included in 100% of prompts to ensure consistent representation. The distributions for age, region of origin, and body type were uniform. The remaining attributes were given heuristic probabilities based on rough estimates of broad, plausible trends (e.g., certain facial hair styles are naturally less common than a clean-shaven look). In the absence of comprehensive demographic data, these estimates were used to illustrate the dataset generation process. Researchers are encouraged to define custom distributions for their specific needs.
Of the 15,000 identities, 111 were excluded due to Arc2Face [40] failing to extract identity embeddings. All images were generated at a resolution of $1024\times 1024$, except for Arc2Face-based live images, which were constrained to $512\times 512$ due to implementation limitations.
@@ 5.2 Similarity-based Identity Filtering
Due to prompt similarity and inherent biases in the FLUX.1 [dev] [3] model, some synthetic identities may appear overly similar. Unlike ONOT [14], which enforces strict identity separation by removing all similar identities detected by a Face Recognition System (FRS), FLUXSynID allows a small degree of similarity, reflecting real-world datasets where a limited False Match Rate (FMR) is acceptable. For instance, FRONTEX guidelines [18] permit an FMR of 0.1%. As the dataset size grows, the probability that a randomly selected identity $i$ falsely matches at least one other identity in the dataset increases, and is defined as:
@@
To manage this behavior, FLUXSynID employs iterative filtering based on FRS similarity scores. Starting from an initial set of document images $D_{o}$, we identify inter-identity pairs that are falsely matched. The filtering process iteratively removes identities involved in the highest number of such matches. This continues until the dataset-wide FMR reaches or falls below a predefined static FMR target:
@@
Experiments in this study leverage two open-source face recognition models: ArcFace [12] and AdaFace [28]. For each, two decision thresholds are selected, corresponding to FMR targets of 0.1% and 0.01%. These thresholds are derived from 340k impostor verification attempts on identities from the CFD dataset, yielding values of 0.423 and 0.497 for ArcFace, and 0.253 and 0.334 for AdaFace, respectively. Accordingly, the $\text{FMR}_{\text{target}}$ in Equation (4) is set to 0.1% or 0.01%, depending on the chosen threshold.
@@
Table 1: Remaining identities after similarity-based filtering at different FMR targets using ArcFace [12] and AdaFace [28]. Percentages are relative to the initial 14,889 identities. Note: lower FMR targets require higher FRS decision thresholds, reducing the number of identity pairs flagged as false matches. Thus, fewer identities are removed and a larger portion of the dataset is retained.
@@
Finally, to ensure there is no identity leakage from LoRA fine-tuning on CFD images, we compared all 14,889 FLUXSynID identities against the 830 CFD subjects used for training. Using AdaFace and a threshold of 0.539, corresponding to the minimum similarity score among mated pairs in the FRGC [41] dataset, no matches were found. This confirms that no identity leakage occurred.
@@ 5.3 Evaluation
Visual Results. Figure 4 shows sample identities from FLUXSynID, each consisting of a document image ($I_{D}$) and live capture images generated by LivePortrait ($I_{LL}$), PuLID ($I_{LP}$), and Arc2Face ($I_{LA}$). Document images are frontal, display neutral expressions, and have uniform white backgrounds, reflecting stylistic control via CFD-based LoRA fine-tuning.
Each live image generation method introduces different levels of variation. LivePortrait introduces subtle, controlled changes in pose and expression while maintaining consistent lighting and background, although minor blurring is often evident. PuLID leverages text, edge maps, and identity embeddings to alter both background and pose, although it may occasionally cause identity drift. Arc2Face generates the most diverse outputs, introducing substantial changes in setting, lighting, and expression that closely mimic in-the-wild photographs.
@@
Prompt-Image Consistency. Figure 5 shows synthetic document images generated from randomly sampled identity attributes. Generally, the images align well with the specified attributes, although some limitations are evident. Age estimation occasionally deviates, producing individuals who appear significantly younger or older than intended, consistent with findings in [39], which highlight FLUX.1 [dev]’s difficulty in accurately representing age.
Figure 5: Examples of randomly sampled identity attributes and their corresponding synthetic document images ($I_{D}$) generated using the FLUXSynID framework.
@@
Empirically, guidance scales between 1.7 and 2.5 yielded the best trade-off between fidelity and prompt consistency. Below 1.7, images often exhibited artifacts and poor prompt adherence, while scales above 2.5 produced overly saturated and visually similar faces, increasing removals during similarity-based filtering.
Impact of Similarity-based Filtering on Attribute Distributions. To evaluate the potential bias introduced by similarity-based filtering (Sec. 5.2), Fig. 6 compares the distributions of gender, age, and region of origin attributes before and after filtering.
Female identities were observed to be removed more frequently, reducing their overall representation. This indicates that FLUX.1 [dev] generates less diverse female faces, leading to higher similarity scores and more frequent exclusion. Conversely, male and non-binary identities increased proportionally, indicating greater variability.
@@
Distribution of Identity Embeddings. Figure 7 visualizes identity embeddings extracted by ArcFace [12] and AdaFace [28] across real and synthetic datasets. Synthetic StyleGAN2 [26] images were generated using the pSp [46] encoder, producing random identities in frontal poses. For ONOT [14], ICAO-compliant images from Subset 1 (the largest identity set) were used. FLUXSynID identities were filtered using FMR of 0.01% thresholds (Sec. 5.2) to remove highly similar identities. Only document images were considered to maintain consistency across datasets, all of which contain document-like image types.
@@
Figure 8: Distribution of cosine similarity scores for mated (blue) and non-mated (orange) identity pairs across real and synthetic datasets, using the AdaFace [28] FRS with an FMR of 0.01% threshold. FLUXSynID-LL, -LP, and -LA refer to live images generated using LivePortrait, PuLID, and Arc2Face, respectively.
@@
Table 2: KL divergence between similarity score distributions of synthetic and real datasets for mated and non-mated identity pairs. Lower values indicate greater distributional similarity. FLUXSynID-LL, -LP, and -LA refer to live images generated using LivePortrait, PuLID, and Arc2Face, respectively.
Similarity Score Distributions. To further evaluate FLUXSynID, cosine similarity scores for mated (document and corresponding live images) and non-mated pairs (document images compared to live images from 100 other identities) are visualized as histograms in Fig. 8, with comparisons to real-world FRLL [11] and FRGC [41] datasets.
@@
These observations are quantified in Tab. 2, which reports KL divergence between synthetic and real datasets. FLUXSynID-LL yields the lowest mated-pair divergence from FRLL, supporting its close visual match and controlled variation. For non-mated pairs, FLUXSynID-LA achieves the lowest divergence, though its mated-pair scores are higher. Overall, FLUXSynID-LL offers the most balanced similarity to FRLL across both pair types.
For FRGC, FLUXSynID-LA records the lowest mated-pair KL divergence, although FLUXSynID-LP visually resembles FRGC images more closely. While its KL values are slightly higher than Arc2Face’s, they still indicate strong distributional similarity.
In contrast, ONOT exhibits high KL divergence across all comparisons, indicating that its similarity score distributions differ substantially from real datasets. This aligns with Fig. 8, where ONOT’s mated and non-mated distributions are clearly shifted relative to real data.
@@ 6 Conclusion
We present FLUXSynID, a synthetic face dataset generation framework capable of producing high-resolution, identity-consistent document and live capture images with control over identity attributes. Combining prompt-driven diffusion synthesis with LoRA fine-tuning [23] and three complementary live image generation methods, FLUXSynID facilitates both controlled and natural variations across identity instances. Experimental results show that FLUXSynID aligns more closely with real-world data distributions than existing synthetic alternatives, while its controlled diversity supports realistic biometric evaluation. These results indicate that FLUXSynID can support biometric tasks requiring structured, identity-consistent image pairs, such as D-MAD, face recognition, and emotion recognition.
Despite its strengths, FLUXSynID has some limitations. Attributes such as age and subtle facial features are not always represented reliably, and similarity-based filtering may unintentionally skew demographic distributions. In contrast to ONOT [14], which enforces strict ICAO compliance [25] at the cost of a 73% reduction in dataset size, FLUXSynID does not perform explicit ICAO filtering. While document-style images are generated to generally align with ICAO standards (e.g., frontal pose, neutral expression), we prioritize dataset size and embedding space diversity over strict adherence, enhancing flexibility for a broader range of biometric research applications.
@@ 7 Acknowledgments
This research was funded by the European Union under the Horizon Europe programme, Grant Agreement No. 101121280. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect the views of the EU/Executive Agency. Neither the EU nor the granting authority can be held responsible for them.